# Rewarding Doubt: A Reinforcement Learning Approach to Calibrated Confidence Expression of Large Language Models

### v3

@@ Abstract
A safe and trustworthy use of Large Language Models (LLMs) requires an accurate expression of confidence in their answers. We propose a novel Reinforcement Learning approach that allows to directly fine-tune LLMs to express calibrated confidence estimates alongside their answers to factual questions. Our method optimizes a reward based on the logarithmic scoring rule, explicitly penalizing both over- and under-confidence. This encourages the model to align its confidence estimates with the actual predictive accuracy. The optimal policy under our reward design would result in perfectly calibrated confidence expressions. Unlike prior approaches that decouple confidence estimation from response generation, our method integrates confidence calibration seamlessly into the generative process of the LLM. Empirically, we demonstrate that models trained with our approach exhibit substantially improved calibration and generalize to unseen tasks without further fine-tuning, suggesting the emergence of general confidence awareness. We provide our training and evaluation code in the supplementary and will make it publicly available upon acceptance.
@@ 1 Introduction
In human intelligence and inter-human interaction, the ability to understand our own uncertainty and communicate our doubts to others is fundamental for effective decision-making, collaboration, and learning [7, 37]. Similarly, for Large Language Models (LLMs) to be safely used in real-world applications, especially when humans and AI systems work together, they must not only generate accurate information but also communicate their confidence in that information. While LLMs have demonstrated impressive capabilities in natural language understanding, question answering and text summarization [32, 6, 1], LLMs still face significant limitations, such as their tendency to generate inaccurate information, often referred to as hallucinations [11]. This raises concerns about their reliability, particularly in real-world applications where trustworthiness is essential. Especially in high-stakes environments such as medical diagnosis, where LLMs are starting to become support tools for professionals [23, 25, 33, 4], overconfident predictions including factual errors or hallucinations could have serious consequences for patient health. Also, in customer service or legal consultation [26, 28], LLMs need to express uncertainty and defer complex queries to human representatives when unsure to avoid misinformed decisions. Reliable confidence estimation and expression would enable these systems to flag uncertain outputs for human review, ensuring that crucial decisions are not made based on uncertain LLM outputs. To allow risk estimation while using LLM-generated output, model confidence should be calibrated, meaning that the expressed numerical confidence should be equal to the probability of the model’s answer being correct.
@@
Many previous methods for confidence estimation lack in calibration performance as they do not train the model and instead infer the confidence from the internal state in a zero-shot setup [14, 19, 8]. Additionally, this does not give models an inherent awareness of confidence. Other trained methods in this area decouple the uncertainty estimation from the text generation process [3, 18]. This approach optimizes for calibrated confidence estimation but does not enable the uncertainty-awareness and expression in the model itself.
Targeting these limitations, we propose a novel reinforcement-learning (RL) approach for teaching LLMs to express their calibrated confidence, encouraging a granular, accurate estimation of the confidence level in the training objective. For this, we model confidence estimation as a betting game: a high-confidence answer would warrant a larger bet, reflecting a strong belief in its correctness, while a lower confidence score would suggest caution. Central to our method is a reward function based on the logarithmic scoring rule, a strictly proper scoring rule. We are the first to optimize this function through reinforcement-learning-based policy optimization, leveraging its calibration properties for directly and seamlessly training confidence calibration in LLM generations. This reward function captures the fundamental risk-reward tradeoff in probabilistic decision-making, as illustrated in Figure 1. It increases the reward when a correct answer is given with high confidence, simulating the higher potential return of big bets. Conversely, it penalizes incorrect answers more when they are made with high confidence, discouraging overconfidence. This ensures that both uncertainty and confidence are appropriately factored into the reward system. As a proper scoring rule, optimizing the reward function trains the model to align its predicted confidence with the accuracy of its output, encouraging granular and calibrated confidence scoring. A calibrated confidence estimation will provably result in the highest reward during training. This not only improves the trustworthiness of LLMs in collaborative human-AI scenarios but also helps users better assess when AI tools should be trusted, double-checked, or deferred to human expertise.
@@ 2 Related Works
Confidence estimation and calibration have a long history in machine learning and natural language processing [34]. With the rise of LLMs, a lot of attention has moved to extracting and improving confidence estimation from these new models. There is a wide range of methods that can roughly be divided into black-box and white-box methods [9].
Black-box methods require only the model generations for uncertainty estimation and do not depend on any knowledge of the model’s internal state. Within this group, linguistic confidence methods assess the uncertainty by simply prompting the LLM to verbalize its confidence. Xiong et al. [37] showed that prompting techniques like chain-of-thought (CoT) [36] can, to a limited extend, help the model in the calibration of its confidence expression. Based on the assumption that uncertainty in its generation would prime the model to hallucinate information, consistency-based methods [21, 35] measure the similarity between multiple generations and show that high variance in the generation correlates with high uncertainty. Wang et al. [35] propose using CoT to produce various reasoning pathways towards an answer. By marginalizing out the reasoning pathways, they provide the most consistent answer as the one most probably correct and offer the consistency measure as an estimation of the model’s confidence. Black-box methods are valuable for their simplicity, ease-of-use and universality, however generally lack behind white-box methods in their calibration performance.
White-box methods, on the other hand, depend on access to the internal state of the model, either for the evaluation of logits and activations or for fine-tuning of the model for calibration. Logit-based methods calibrate the confidence by assessing the probabilities or the entropy of tokens in the predicted sequence. Huang et al. [14] calculate the mean probability of the generated tokens with the intuition that a model that generates an answer with high probability tokens also has a high confidence in this answer. Kuhn et al. [19] and [8] expand on this idea by looking at the entropy of the generated tokens while also taking the semantic relevance into account. Kadavath et al. [17] showed that LLMs can self-evaluate the correctness of their own generation. They prompt the model to provide an answer followed by a judgement whether its own answer is "true" or "false". They then compare the probability of the "true" or "false" token to calculate a confidence estimation.
While all the aforementioned methods are zero-shot methods, another line of white-box methods, proposes to train external models for confidence estimation. Xiong et al. [37] propose a method that trains a linear probe for assessing the model’s confidence based on the activations of the LLM’s hidden state. They evaluate the application of the probe to various layers and train it with correctness labels of the task. This approach does not explicitly train on the granularity of confidence estimation and, therefore, is only indirectly able to express a confidence gradient for fine-grained calibration. Building on this approach, Mielke et al. [22] give an LLM access to the confidence extracted by the external probe and fine-tune it to express this confidence. However, the calibration performance is upper-bounded by the ability of the trained probe to assess the model’s confidence accurately. Stengel-Eskin et al. [27] propose a DPO-based method, LACIE, simulating an interaction between a generation and a listener model, where the speaker model generates answers, and the listener model evaluates them based on perceived confidence. By training with Direct Preference Optimization (DPO) on data from TriviaQA, LACIE rewards accurate confidence expression, including correct answers accepted by the listener model and incorrect, rejected ones, while penalizing misleading confidences. The confidence expression is calibrated to how the listener model interprets confidence cues, depending on its ability to judge confidence and aligning perceived confidence rather than a fact-based numerical confidence score. However, none of these methods provide theoretical guarantees or provable optimality regarding the trained calibration of the expressed confidence.
Simultaneously, fine-tuning LLMs using Reinforcement Learning from human feedback (RLHF) has received increasing attention for aligning the models with human preferences, avoiding toxicity or the generation of harmful content [24, 20]. Outside of RLHF, RL has been used to fine-tune LLMs for agentic interaction in textual environments [40, 5]. Confidence estimation of LLMs using RL is a novel research direction that, to the best of our knowledge, has only been approached by Tao et al. [30] and Xu et al. [38]. Tao et al. [30] build on the Reinforcement Learning with Human Feedback (RLHF) paradigm and design a reward function that aligns the verbalized confidence of the model with the preference rating predicted by the RLHF reward model. However, this approach can not evaluate model confidence in factual statements and depends on human-annotated preference data. Xu et al. [38] introduce a novel Reinforcement Learning from Knowledge Feedback (RLKF) framework, which aims to teach a model to refuse to answer questions outside their knowledge scope, reducing hallucinations, but does not quantify confidence when giving an answer.
@@
In contrast, our method directly optimizes for factual calibration using a theoretically grounded, proper scoring rule as the reward signal, enabling the model to develop intrinsic uncertainty awareness without requiring external preference models, knowledge supervision, or post-hoc calibration techniques, while at the same time seamlessly integrating calibrated confidence expression into the LLMs response generation.
@@ 3 Rewarding Doubt
We propose a novel reinforcement-learning approach, that improves an LLM’s ability to verbalize an accurate numerical confidence in a previously generated answer. The LLM functions as an agent in a simulated environment as shown in Figure 2, that poses challenging question-answering scenarios. It is prompted with task queries such as factual questions and asked to predict both an answer to the query as well as a confidence score. Based on the correctness of the answer, and the expressed confidence, we reward the model, incentivizing it to express a calibrated confidence.
@@
Formally, let the model be provided with a textual question or request $q$, resulting in an answer-confidence pair $(a,\hat{p})$ as response, where $a$ is a textual answer with binary correctness value, and $0\leq\hat{p}\leq 1$ is a numerical confidence score representing the subjective probability the model assigns to answer $a$ being correct. We train this subjective probability assessment to align with the true epistemic probability $p^{*}$, which represents the actual likelihood of correctness given the model’s internal knowledge state. If $\hat{p}$ and $p^{*}$ are aligned the model is perfectly calibrated, meaning the probability of correctness $P(j(a)=1)$ always equals the expressed confidence:
@@
To promote accurate confidence estimation, the model’s reward function must fulfill the requirement of being maximized when $\hat{p}=p^{*}$. The model should receive a higher reward when it correctly predicts an outcome with high confidence. Conversely, incorrect predictions should be penalized more severely when the model exhibits high confidence. In contrast, expressing uncertainty reduces potential penalties, leading to a higher expected reward in uncertain scenarios. This approach incentivizes the model to express high confidence only in cases where certainty is warranted while expressing doubt in ambiguous situations. By penalizing both overconfidence and underconfidence, the model is encouraged to calibrate its confidence accurately, effectively balancing the trade-off between reward maximization and penalty avoidance.
We design our reward as a logarithmic scoring function, fulfilling the described optimality conditions:
$R(a,\hat{p},j)=\begin{cases}log(\max(\hat{p},\epsilon)),&\text{if $j(a)$ = 1 (correct)}\\ log(\min(1-\hat{p},1-\epsilon)),&\text{\text{if $j(a)$ = 0 (incorrect)}}\end{cases}$
where $\epsilon>0$ is a small positive constant introduced for numerical stability to avoid evaluating the logarithm at zero. The normalized reward for correct and incorrect answers for each confidence is visualized in Figure 3.
@@
The proof of Proposition 1 is analogous to the proof that the logarithmic scoring rule is a proper scoring rule. We provide it in full in Appendix B and discuss the influence of the clipping on the optimality of the reward function.
@@ 4 Experimental Setup
We evaluate our method in both Single-Answer and Multiple-Answer settings. We prompt the model to provide a confidence for each answer as an integer between 0 and 10, which we normalize for the reward calculation. A confidence of zero is defined as the model being certain that the answer is incorrect, while ten is defined as the model being certain the answer is correct. We normalize the reward function to the range of $[-1,1]$.
In the Single-Answer setting we train the model on the TriviaQA dataset [16], which contains question-answer-evidence triplets, from which we only use the questions and answers. For generalization experiments, we evaluate our method on CommonsenseQA [29] and MedQA [15], which are multiple-choice question datasets in the commonsense and medical domain, respectively. For the Multiple-Answer setting, we train on the QAMPARI dataset [2], which contains questions with multiple-answers as well as evidence, again only using the questions and answers.
In the Single-Answer setting we compare our approach on the TriviaQA dataset against the following methods: Chain-of-Thought [37], Top-K [31], Surrogate Token [17], Sequence Probability [14] and Self-Consistency [35] as zero-shot methods, LACIE [27], which uses DPO for optimizing confidence expression and Trained Probe [3], which employs supervised training of an external probe for estimation model. We also compare to the non-finetuned base model in a zero-shot manner, using the same prompt as our Rewarding Doubt method and refer to this setup as Verbalize. In the Multiple-Answer setting we compare to Trained Probe and Sequence probability, as those methods are the best performing zero-shot and trained baselines in the Single-Answer setting. LACIE does not report results for this dataset, thus we can only compare on TriviaQA.
We report our results using the Expected Calibration Error (ECE) and the Area Under the Receiver Operating Characteristic Curve (AUROC) metric. The ECE measures the difference between a model’s predicted confidence and its actual accuracy, assessing how well confidences align with real outcomes. AUROC assesses a model’s ability to distinguish between classes, with a score of 0.5 indicating random guessing and 1.0 indicating perfect classification. Additionally, we visualize the calibration with calibration curves, where a well-calibrated model lies close to the 45° line and large deviations show a high miscalibration.
@@
For the multiple-choice datasets MedQA and CommonsenseQA, we evaluate correctness using the exact string matching between the model’s response and the ground truth answer.
For the TriviaQA and QAMPARI datasets, we use the F1 score of word overlap to measure the similarity between the model’s response and the ground truth candidates. The F1 score is calculated for each candidate and the maximum score is considered the final score. We consider an answer as correct if its score exceeds a threshold of 0.5.
5 Implementation Details
Training
As base model for our experiments we use Meta-Llama-3-8B-Instruct [10]. We employ the 4-bit quantized performance-optimized model version by Unsloth AI [12]. Additionally, we apply LoRA fine-tuning [13], reducing the amount of trainable parameters from 8,030,265,345 to 41,947,137. For the Single-Answer setting we train the model for two epochs with a learning rate of 1e-5. For the Multiple-Answer setting, due to the size of the training dataset and the fact that each question yields multiple facts, the model is trained for a limited amount of 24,000 steps with a batchsize of eight and a learning rate of 1e-5 with a reward scale of five. All models are trained on one Nvidia A40 with each training run taking seven days. On average the model generated approximately 3.4 answers per fact. If the model fails to generate an answer in the specified format, it is penalized with an out-of-format reward of -3.
Baselines
For the Sequence Probability, we compute the average probability for each token in the response. In the Self-Consistency method, we let the model explore ten reasoning pathways, and the similarity of each resulting output is evaluated using the BERTScore metric [39]. For the trained probe [3], the original study introduced a custom dataset comprising short statements classified as either true or false. The model’s activations in response to these statements were extracted from specific layers, and a multilayer perceptron (MLP) was subsequently trained on these activations to predict the truthfulness of the statements. To ensure a fair comparison, we adapted this methodology to better align with our data by allowing the model to generate answers to training dataset questions and then extracting its activations from the 24th layer for both the statements and their corresponding answers. The labels for each sample were determined following the same evaluation procedure as described in our evaluation framework. For the architecture of the MLP, we employed the same design as Azaria and Mitchell [3] and train it for four epochs with a learning rate of 1e-4 until convergence. The exact prompts used for each baseline are provided in Appendix A.

### v4

@@ Abstract
A safe and trustworthy use of Large Language Models (LLMs) requires an accurate expression of confidence in their answers. We propose a novel Reinforcement Learning approach that allows to directly fine-tune LLMs to express calibrated confidence estimates alongside their answers to factual questions. Our method optimizes a reward based on the logarithmic scoring rule, explicitly penalizing both over- and under-confidence. This encourages the model to align its confidence estimates with the actual predictive accuracy. The optimal policy under our reward design would result in perfectly calibrated confidence expressions. Unlike prior approaches that decouple confidence estimation from response generation, our method integrates confidence calibration seamlessly into the generative process of the LLM. Empirically, we demonstrate that models trained with our approach exhibit substantially improved calibration and generalize to unseen tasks without further fine-tuning, suggesting the emergence of general confidence awareness. Our code is available at https://github.com/pasta99/RewardingDoubt.
@@ 1 Introduction
In human intelligence and inter-human interaction, the ability to understand our own uncertainty and communicate our doubts to others is fundamental for effective decision-making, collaboration, and learning (Cosmides & Tooby, 1996; Xiong et al., 2024). Similarly, for Large Language Models (LLMs) to be safely used in real-world applications, especially when humans and AI systems work together, they must not only generate accurate information but also communicate their confidence in that information. While LLMs have demonstrated impressive capabilities in natural language understanding, question answering and text summarization (Touvron et al., 2023; Chiang et al., 2023; Achiam et al., 2023), LLMs still face significant limitations, such as their tendency to generate inaccurate information, often referred to as hallucinations (Hadi et al., 2023). This raises concerns about their reliability, particularly in real-world applications where trustworthiness is essential. Especially in high-stakes environments such as medical diagnosis, where LLMs are starting to become support tools for professionals (Moor et al., 2023; Pellegrini et al., 2025; Tu et al., 2024; Bani-Harouni et al., 2024), overconfident predictions including factual errors or hallucinations could have serious consequences for patient health. Also, in customer service or legal consultation (Shi et al., 2024; Sun et al., 2024), LLMs need to express uncertainty and defer complex queries to human representatives when unsure to avoid misinformed decisions. Reliable confidence estimation and expression would enable these systems to flag uncertain outputs for human review, ensuring that crucial decisions are not made based on uncertain LLM outputs. To allow risk estimation while using LLM-generated output, model confidence should be calibrated, meaning that the expressed numerical confidence should be equal to the probability of the model’s answer being correct.
@@
Many previous methods for confidence estimation lack in calibration performance as they do not train the model and instead infer the confidence from the internal state in a zero-shot setup (Huang et al., 2023; Kuhn et al., 2023; Duan et al., 2024). Additionally, this does not give models an inherent awareness of confidence. Other trained methods in this area decouple the uncertainty estimation from the text generation process (Azaria & Mitchell, 2023; Kapoor et al., 2024). This approach optimizes for calibrated confidence estimation but does not enable the uncertainty-awareness and expression in the model itself.
Targeting these limitations, we propose a novel reinforcement-learning (RL) approach for teaching LLMs to express their calibrated confidence, encouraging a granular, accurate estimation of the confidence level in the training objective. For this, we model confidence estimation as a betting game: a high-confidence answer would warrant a larger bet, reflecting a strong belief in its correctness, while a lower confidence score would suggest caution. Central to our method is a reward function based on the logarithmic scoring rule, a strictly proper scoring rule. We are the first to optimize this function through reinforcement-learning-based policy optimization, leveraging its calibration properties for directly and seamlessly training confidence calibration in LLM generations. This reward function captures the fundamental risk-reward tradeoff in probabilistic decision-making, as illustrated in Figure˜1. It increases the reward when a correct answer is given with high confidence, simulating the higher potential return of big bets. Conversely, it penalizes incorrect answers more when they are made with high confidence, discouraging overconfidence. This ensures that both uncertainty and confidence are appropriately factored into the reward system. As a proper scoring rule, optimizing the reward function trains the model to align its predicted confidence with the accuracy of its output, encouraging granular and calibrated confidence scoring. A calibrated confidence estimation will provably result in the highest reward during training. This not only improves the trustworthiness of LLMs in collaborative human-AI scenarios but also helps users better assess when AI tools should be trusted, double-checked, or deferred to human expertise.
@@ 2.1 Confidence Estimation in LLMs
Confidence estimation and calibration have a long history in machine learning and natural language processing (Wang, 2024). With the rise of LLMs, research has focused on adapting and extending these ideas to modern architectures. Broadly, methods fall into black-box and white-box approaches (Geng et al., 2024).
Black-box methods
Black-box methods rely only on model outputs. Linguistic prompting methods ask the model to verbalize its confidence, sometimes aided by chain-of-thought reasoning (Xiong et al., 2024; Wei et al., 2022). Consistency-based approaches estimate confidence by measuring agreement across multiple generations, with high variance indicating uncertainty (Manakul et al., 2023; Wang et al., 2022). Recently, Zhou et al. (2025) proposed SteerConf, which does multiple inference passes where the LLM is prompted to use different levels of caution in its confidence expression. The resulting verbalized confidences are aggregated based on confidence and answer consistency to an overall confidence prediction. Black-box methods are valuable for their simplicity, ease-of-use and universality, however generally lack behind white-box methods in their calibration performance.
White-box methods
White-box methods exploit internal model states. Logit-based techniques estimate confidence from token probabilities or entropy (Huang et al., 2023; Kuhn et al., 2023; Duan et al., 2024), assuming that high probability tokens correspond to high confidence predictions. Self-evaluation methods let the model judge the truth of its own answers (Kadavath et al., 2022). They prompt the model to provide an answer followed by a judgment whether its own answer is "true" or "false". They then compare the probability of the "true" or "false" token to calculate a confidence estimation. External probing approaches train classifiers on hidden states to predict correctness (Azaria & Mitchell, 2023). While some of these methods achieve good confidence estimation results, they do not teach the model to express clear confidence values itself but depend on some auxiliary estimation mechanism.
2.2 Finetuned Confidence Expression
A growing line of work integrates confidence estimation into instruction tuning. These methods typically follow a two-step paradigm: First, they estimate model confidence using various methods, e.g., self-consistency (Cheng et al., 2024; Yang et al., 2024; Han et al., 2024), token probabilities (Chen et al., 2024), trained probes (Mielke et al., 2022), empirical accuracy (Zhang et al., 2024; Lin et al., 2022; Ulmer et al., 2024), or topic unfamiliarity (Wan et al., 2024; Kang et al., 2024). Second, they construct finetuning datasets that either replace uncertain answers with refusals (Zhang et al., 2024; Cheng et al., 2024; Yang et al., 2024; Wan et al., 2024) or append the estimated uncertainty as an additional supervised signal (Han et al., 2024; Chen et al., 2024; Mielke et al., 2022; Lin et al., 2022; Ulmer et al., 2024).
The key limitation of this approach is that the model’s expressed confidence is bounded by the quality of the constructed ground-truth estimates. Additionally, while the underlying confidence estimation method might optimize for perfect calibration (e.g. in the case of the trained probe), this theoretical guarantee is lost when performing supervised finetuning on these constructed ground truths to reproduce these scores.
2.3 Reinforcement Learning for Confidence Expression
Reinforcement Learning from Human Feedback (RLHF) has proven effective for aligning LLMs with human preferences (Ouyang et al., 2022), and has also been explored for agentic interaction in textual environments (Zhou et al., 2023; Carta et al., 2023). Only recently have researchers begun applying RL directly to confidence estimation. Tao et al. (2024) adapt RLHF by designing rewards that align verbalized confidence with preference ratings, but this requires human-annotated preference data and does not address factual calibration. Leng et al. (2024) identify that standard reward models in RLHF are biased toward high verbalized confidence, rating answers with high confidence expressions with a high reward. To counteract this, they introduce two reward model training paradigms, PPO-M and PPO-C, which fine-tune the reward model to reward answers where correctness and confidence expression are aligned. Xu et al. (2024) propose RL from Knowledge Feedback (RLKF) to encourage refusals outside the model’s knowledge scope, reducing hallucinations but without quantifying confidence. Stengel-Eskin et al. (2024) propose LACIE, a DPO-based approach that simulates an interaction between a speaker and a listener model, rewarding accurate and honest confidence expression by aligning it with the listener’s interpretation of confidence cues rather than with fact-based numerical calibration.
@@ 2.3 Reinforcement Learning for Confidence Expression
In contrast to previous works, our method directly optimizes for factual calibration using a theoretically grounded, proper scoring rule as the reward signal, enabling the model to develop intrinsic uncertainty awareness without requiring external preference models, knowledge supervision, or post-hoc calibration techniques, while at the same time seamlessly integrating calibrated confidence expression into the LLMs response generation.
@@ 3 Rewarding Doubt
We propose a novel reinforcement-learning approach, that improves an LLM’s ability to verbalize an accurate numerical confidence in a previously generated answer. The LLM functions as an agent in a simulated environment as shown in Figure˜2, that poses challenging question-answering scenarios. It is prompted with task queries such as factual questions and asked to predict both an answer to the query as well as a confidence score. Based on the correctness of the answer, and the expressed confidence, we reward the model, incentivizing it to express a calibrated confidence.
@@
[…]
@@
To promote accurate confidence estimation, the model’s expected reward must fulfill the requirement of being maximized when $\hat{p}=p^{*}$, i.e. when the predicted confidence aligns with the probability of correctness. The model should receive a high reward when it correctly predicts an outcome with high confidence or when it incorrectly predicts an outcome with low confidence. Conversely, the reward should be low when incorrect predictions are provided with high confidence or correct predictions are provided with low confidence. This approach incentivizes the model to express high confidence only in cases where certainty is warranted while expressing doubt in ambiguous situations. By penalizing both overconfidence and underconfidence, the model is encouraged to calibrate its confidence accurately, effectively balancing the trade-off between reward maximization and penalty avoidance. Note, that through this design our method focuses exclusively on improving calibration while keeping task performance stable.
We design our reward as a logarithmic scoring function:
[…]
This function fulfills the requirement described above as we show in the following proposition:
@@
The proof of Proposition 1 is analogous to the proof that the logarithmic scoring rule is a proper scoring rule. We provide it in full in Appendix˜B and discuss the influence of the clipping on the optimality of the reward function.
Since the logarithm of zero is undefined, we introduce a small positive constant $\epsilon$ as clipping value for numerical stability. Concretely, we clip the lower and upper limit of the confidence $\hat{p}$ to $\epsilon$ and $1-\epsilon$, respectively. The clipped reward function is provided in Appendix˜C. The normalized and clipped reward for correct and incorrect answers for each confidence is visualized in Figure˜3.
@@ 4 Experimental Setup
We evaluate our method in both Single-Answer and Multiple-Answer settings. We prompt the model to provide a confidence for each answer as an integer between 0 and 10, which we normalize for the reward calculation. A confidence of zero is defined as the model being certain that the answer is incorrect, while ten is defined as the model being certain the answer is correct. We normalize the reward function to the range of [1,1][-1,1].
In the Single-Answer setting we train the model on the TriviaQA dataset (Joshi et al., 2017), which contains question-answer-evidence triplets, from which we only use the questions and answers. For generalization experiments, we evaluate our method on CommonsenseQA (Talmor et al., 2019) and MedQA (Jin et al., 2020), which are multiple-choice question datasets in the commonsense and medical domain, respectively. For the Multiple-Answer setting, we train on the QAMPARI dataset (Amouyal et al., 2023), which contains questions with multiple-answers as well as evidence, again only using the questions and answers.
In the Single-Answer setting we compare our approach on the TriviaQA dataset against the following methods: Chain-of-Thought (Xiong et al., 2024), Top-K (Tian et al., 2023), Surrogate Token (Kadavath et al., 2022), Sequence Probability (Huang et al., 2023) and Self-Consistency (Wang et al., 2022) as zero-shot methods, LACIE (Stengel-Eskin et al., 2024), which uses DPO for optimizing confidence expression and Trained Probe (Azaria & Mitchell, 2023), which employs supervised training of an external probe for estimation model. We also compare to the non-finetuned base model in a zero-shot manner, using the same prompt as our Rewarding Doubt method and refer to this setup as Verbalize. In the Multiple-Answer setting we compare to Trained Probe and Sequence probability, as those methods are the best performing zero-shot and trained baselines in the Single-Answer setting. LACIE does not report results for this dataset, thus we can only compare on TriviaQA.
We report our results using the Expected Calibration Error (ECE) and the Area Under the Receiver Operating Characteristic Curve (AUROC) metric. Additionally, we visualize the calibration with calibration curves, where a well-calibrated model lies close to the 45° line and large deviations show a high miscalibration.
@@
For the multiple-choice datasets MedQA and CommonsenseQA, we evaluate correctness using the exact string matching between the model’s response and the ground truth answer. For the TriviaQA and QAMPARI datasets, we use the F1 score of word overlap to measure the similarity between the model’s response and the ground truth candidates. The F1 score is calculated for each candidate and the maximum score is considered the final score. We consider an answer as correct if its score exceeds a threshold of 0.5.
Implementation Details
We optimize the reward function using the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017). Unless stated otherwise, we use Meta-Llama-3-8B-Instruct (Grattafiori et al., 2024) as base model for our experiments. We employ the 4-bit quantized performance-optimized model version by Unsloth AI (Han et al., 2023) and apply LoRA fine-tuning (Hu et al., 2022). For the Single-Answer setting we train the model for two epochs with a learning rate of 1e-5. For the Multiple-Answer setting, due to the size of the training dataset and the fact that each question yields multiple facts, the model is trained for a limited amount of 24,000 steps with a batchsize of eight and a learning rate of 1e-5 and multiply the reward with 5 to increase its spread. All models are trained on one Nvidia A40 with each training run taking seven days. On average the model generated approximately 3.4 answers per fact. If the model fails to generate an answer in the specified format, it is penalized with an out-of-format reward of -3. Detailed implementation choices for the baselines are provided in Appendix˜D.